OpenStudy (ohmath123):

7•4^4x-5=72 How can I solve this using the rules of logarithms?

Nnesha (nnesha):

\[\large \rm 7 \cdot 4^{4x-5}=72 ~~~OR~~~~7\cdot 4^{4x}-5 =72 \] which one ? i think 2nd one

OpenStudy (ohmath123):

the first one @Nnesha

Nnesha (nnesha):

alright. in order to drop the power first we need to get rid of any number/ variable front of the base so what would you do to cancel out 7 from left side ?

OpenStudy (ohmath123):

log base of 7? @Nnesha

OpenStudy (ohmath123):

to both sides?

Nnesha (nnesha):

hmm we should cancel out 7 from left side fist than deal with log it's 7 times 4^{4x-5} to cancel out 7 divide both sides by 7

OpenStudy (ohmath123):

so im left with 4^4x-5=72/7 @Nnesha

Nnesha (nnesha):

right now you can take log both sides.

OpenStudy (ohmath123):

so that would be log 4^4x-5= log 72/7 . Then i can bring the 4x-5 to the front like this: 4x-5 log 4=log 72/7? @Nnesha

Nnesha (nnesha):

looks good

Nnesha (nnesha):

i would divide 72/7 to get decimal

OpenStudy (ohmath123):

then divide log 4

OpenStudy (ohmath123):

Thank you! i got it from here. Thank you so much for guiding me! @Nnesha

Nnesha (nnesha):

alright good :)

Nnesha (nnesha):

my pleasure :)